The following balances appeared in the records of a company on 28 February 2012, the end of the financial year: Vehicles 330 000 Accumulated depreciation: Vehicles 75 000 INSTRUCTIONS: Complete the extract from the PPE Note at the end of the financial year, 28 February 2012, after taking the following transactions in consideration: 1. On 30 Nov 2011 a vehicle, bought on 1 July 2009 for R110 000, was sold on credit, for R55 000. Depreciation was provided for on the diminishing balance method at 25% p.a. on all vehicles, unless otherwise stipulated. 2. A new delivery vehicle, with an estimated life of 4 years and a residual value of R15 000, was purchased on 1 Sept 2011 for R135 000. An additional R5000 was paid for a new sound system. The straight line method of depreciation must be used for this delivery vehicle only.


1
Expert's answer
2021-10-19T10:02:53-0400

Solution:

1.). MV cost = 110,000

Calculate depreciation using reducing balance method:

Depreciation 2009 = 110,000 "\times 25\%" "\times" 6/12 = 13,750

Depreciation 2010 = 96,250 "\times 25\%" = 24,063

Depreciation 2011 = 72,817 "\times 25\%" "\times" 11/12 = 16,543

Total depreciation = 13,750 + 24,063 + 16,543 = 54,356

Dr. Depreciation A/c                                             54,356

             Cr. Accumulated depreciation A/c                                     54,356

 

Loss on disposal = 72,187 – 16,543 – 55,000 = 644

Dr. Cash A/c                                                          55,000

Dr. Accumulated Depreciation A/c                       54,356

Dr. Loss on disposal Vehicle A/c                               644

                      Cr. Vehicle A/c                                                           110,000

 

2.). Purchase of Vehicle:

Dr. Vehicle A/c                                               135,000

                   Cr. Cash A/c                                                                   135,000

Depreciation: 135,000 – 15,000/4 = 30,000

Dr. Depreciation                                                 30,000

                  Cr. Accumulated Depreciation A/c                                      30,000

 

Vehicle and accumulated depreciation balances as at February 28th 2012:

New vehicle cost = 330,000 + 135,000 – 54,356 – 30,000 – 644 = 380,000

New accumulated depreciation vehicle = 75,000 + 54,356 + 30,000 + 644 = 160,000